# Watching the Watchers: Exposing Gender Disparities in Machine Translation Quality Estimation

### v3

@@ Abstract
Quality estimation (QE)—the automatic assessment of translation quality—has recently become crucial across several stages of the translation pipeline, from data curation to training and decoding. While QE metrics have been optimized to align with human judgments, whether they encode social biases has been largely overlooked. Biased QE risks favoring certain demographic groups over others, e.g., by exacerbating gaps in visibility and usability. This paper defines and investigates gender bias of QE metrics and discusses its downstream implications for machine translation (MT). Experiments with state-of-the-art QE metrics across multiple domains, datasets, and languages reveal significant bias. When a human entity’s gender in the source is undisclosed, masculine-inflected translations score higher than feminine-inflected ones and gender-neutral translations are penalized. Even when contextual cues disambiguate gender, using context-aware QE metrics leads to more errors in picking the correct translation inflection for feminine than masculine referents. Moreover, a biased QE metric affects data filtering and quality-aware decoding. Our findings highlight the need for renewed focus in developing and evaluating QE metrics centered around gender.111We release code and artifacts at: https://github.com/g8a9/gender-bias-qe-metrics.
@@ 1 Introduction
In this paper, we conceptualize and measure gender bias in QE metrics and its downstream effects. We draw two conditions to identify a biased metric. A QE metric exhibits gender bias if, when translating human entities, it systematically (i) assigns higher scores to one gender form over others when no explicit lexical cues disclose gender or (ii) fails to detect the correct gender form even when lexical cues are provided, disproportionately across genders. We investigate bias across commonly used QE metrics through an extensive empirical study. We focus on out-of-English translation where the target language uses grammatical gender,222I.e., where gender for human referents is assigned on a semantic basis and reflected with morphological marks. a scenario frequently used in prior research on gender bias in MT (e.g., Vanmassenhove et al., 2018; Stanovsky et al., 2019; Saunders et al., 2020; Attanasio et al., 2023; Sant et al., 2024). We address both binary (i.e., contrasting masculine and feminine translations) and gender-neutral (i.e., contrasting gendered and gender-neutral forms) translation scenarios, and experiment with 11 state-of-the-art metrics (supervised and prompted) across two domains, three datasets, and eight target languages.
@@ 2 Related Work
Evaluating MT metrics
Significant efforts have been made to evaluate automatic metrics for MT. Most have focused on comparing sentence- or system-level correlations between a metric’s score and human judgements (Macháček and Bojar, 2014; Stanojević et al., 2015; Kocmi et al., 2021; Freitag et al., 2023; Deutsch et al., 2023; Thompson et al., 2024). Finer-grained analyses have discovered poor robustness to low-quality translations (Fomicheva and Specia, 2019), domain (Zouhar et al., 2024), or passive voice (Avramidis et al., 2023). Metrics have also been shown to produce imbalanced score distributions (Sun et al., 2020), and fail to distinguish discrepancies in numbers and named entities (Amrhein and Sennrich, 2022) or high-quality translations (Agrawal et al., 2024). However, none of the studies have thus far evaluated gender bias exhibited by QE metrics.
Social biases in NLG metrics
A few studies have assessed whether metrics used to evaluate natural language generation (NLG) exhibit social biases. Qiu et al. (2023) test whether n-gram- and model-based metrics show gender bias in image captioning. Sun et al. (2022) quantify several social biases. They use English coreference resolution examples to measure the gender bias of several n-grams and model-based NLG metrics. Gao and Wan (2022) measure race and gender stereotypes using WEAT (Caliskan et al., 2017) and SEAT (May et al., 2019). However, no study targets QE metrics for MT or reference-free NLG metrics.
@@ 4 Findings
4.1 Gender-Ambiguous Entities
@@ 6 Conclusion
Our study is characterized by several methodological constraints that merit critical reflection. We limited our investigation to sentence-level MT, which, while representative of prevalent machine translation scenarios, does not cover phenomena potentially arising within more extended contexts, e.g., document-level Jiang et al. (2022); Raunak et al. (2023) and conversational MT Agrawal et al. (2024). The English source sentences examined were derived from naturalistic passages of high quality and relatively limited syntactic variation.
Moreover, our evaluation of neural QE metrics with extra-sentential context, relies on inference-time strategies for integrating contextual information Vernikos et al. (2022); Agrawal et al. (2024). We acknowledge that translating context may introduce additional bias depending on the specific translation system used. We leave a more in-depth investigation of this phenomenon, including more resource-intensive solutions—such as targeted fine-tuning—to future work.

### v4

@@ Abstract
Quality estimation (QE)—the automatic assessment of translation quality—has recently become crucial across several stages of the translation pipeline, from data curation to training and decoding. While QE metrics have been optimized to align with human judgments, whether they encode social biases has been largely overlooked. Biased QE risks favoring certain demographic groups over others, e.g., by exacerbating gaps in visibility and usability. This paper defines and investigates gender bias of QE metrics and discusses its downstream implications for machine translation (MT). Experiments with state-of-the-art QE metrics across multiple domains, datasets, and languages reveal significant bias. When a human entity’s gender in the source is undisclosed, masculine-inflected translations score higher than feminine-inflected ones, and gender-neutral translations are penalized. Even when contextual cues disambiguate gender, using context-aware QE metrics leads to more errors in selecting the correct translation inflection for feminine referents than for masculine ones. Moreover, a biased QE metric affects data filtering and quality-aware decoding. Our findings underscore the need for a renewed focus on developing and evaluating QE metrics centered on gender.111We release code and artifacts at: https://github.com/deep-spin/gender-bias-qe-metrics.
@@ 1 Introduction
In this paper, we conceptualize and measure gender bias in QE metrics and its downstream effects. We draw two conditions to identify a biased metric. A QE metric exhibits gender bias if, when translating human entities, it systematically (i) assigns higher scores to one gender form over others when no explicit lexical cues disclose gender, or (ii) fails to detect the correct gender form even when lexical cues are provided, disproportionately across genders. We investigate bias across commonly used QE metrics through an extensive empirical study. We focus on out-of-English translation where the target language uses grammatical gender,222I.e., where gender for human referents is assigned on a semantic basis and reflected with morphological marks. a scenario frequently used in prior research on gender bias in MT (e.g., Vanmassenhove et al., 2018; Stanovsky et al., 2019; Saunders et al., 2020; Attanasio et al., 2023; Sant et al., 2024). We address both binary (i.e., contrasting masculine and feminine translations) and gender-neutral (i.e., contrasting gendered and gender-neutral forms) translation scenarios, and experiment with 11 state-of-the-art metrics (supervised and prompted) across two domains, three datasets, and eight target languages.
@@ 2 Related Work
2.1 Evaluating MT Metrics
Significant efforts have been made to evaluate automatic metrics for MT. Most have focused on comparing sentence- or system-level correlations between a metric’s score and human judgements (Macháček and Bojar, 2014; Stanojević et al., 2015; Kocmi et al., 2021; Freitag et al., 2023; Deutsch et al., 2023; Thompson et al., 2024). Finer-grained analyses have discovered poor robustness to low-quality translations (Fomicheva and Specia, 2019), domain (Zouhar et al., 2024), or passive voice (Avramidis et al., 2023). Metrics have also been shown to produce imbalanced score distributions (Sun et al., 2020), and fail to distinguish discrepancies in numbers and named entities (Amrhein and Sennrich, 2022) or high-quality translations (Agrawal et al., 2024b). However, none of the studies have thus far evaluated gender bias exhibited by QE metrics.
2.2 Gender Bias in MT
Gender bias in MT systems—stemming from societal norms, model design, and deployment decisions—has been extensively documented (Savoldi et al., 2021). Prior work in NLP has primarily focused on measuring the extent of such bias in system outputs (e.g., Stanovsky et al., 2019; Cho et al., 2019; Hovy et al., 2020; Vanmassenhove and Monti, 2021; Levy et al., 2021; Lucy and Bamman, 2021; Sant et al., 2024; Piergentili et al., 2023; Lauscher et al., 2023; Lardelli et al., 2024; Stewart and Mihalcea, 2024), as well as on proposing mitigation strategies (e.g., Escudé Font and Costa-jussà, 2019; Costa-jussà and de Jorge, 2020; Saunders et al., 2020; Saunders and Byrne, 2020; Escolano et al., 2021; Lee et al., 2023; Garg et al., 2024). In contrast to this line of work, we shift the focus from the translation systems themselves to the task of quality estimation, a step that, typically, follows the MT. Further, we investigate whether bias in QE can propagate to or reinforce bias in MT outputs.
2.3 Social Biases in NLG Metrics
Several studies have investigated whether metrics used to evaluate natural language generation (NLG) exhibit social biases. Qiu et al. (2023) test whether n-gram- and model-based metrics, e.g. CLIPScore (Hessel et al., 2021) show gender bias in image captioning. Sun et al. (2022) quantify several social biases. They use English coreference resolution examples to measure the gender bias of several n-grams and model-based NLG metrics, including BERTScore (Zhang* et al., 2020), MoverScore (Zhao et al., 2019), and BLEURT (Sellam et al., 2020). Gao and Wan (2022) measure race and gender stereotypes using WEAT (Caliskan et al., 2017) and SEAT (May et al., 2019). However, no study targets QE metrics for MT or reference-free NLG metrics.
@@ 4 Findings
4.1 Gender-Ambiguous Instances
@@ 6 Conclusion
Our study is characterized by several methodological constraints that merit critical reflection. We limited our investigation to sentence-level MT, which, while representative of prevalent machine translation scenarios, does not cover phenomena potentially arising within more extended contexts, e.g., document-level Jiang et al. (2022); Raunak et al. (2023); Fernandes et al. (2025) and conversational MT Agrawal et al. (2024a). The English source sentences examined were derived from naturalistic passages of high quality and relatively limited syntactic variation.
Moreover, our evaluation of neural QE metrics with extra-sentential context, relies on inference-time strategies for integrating contextual information Vernikos et al. (2022); Agrawal et al. (2024a). We acknowledge that translating context may introduce additional bias depending on the specific translation system used. We leave a more in-depth investigation of this phenomenon—including more resource-intensive solutions, such as targeted fine-tuning or interpretability-driven strategies for selecting few-shot examples to mitigate biases (e.g., gender) in the translation of contextual information Thakur et al. (2023); Zaranis et al. (2024)—to future work.
@@
Acknowledgements
We would like to thank Ben Peters, Dennis Fucci, António Farinhas, Sonal Sannigrahi for their feedback and insightful discussions in the earlier versions of the paper, and Anna Currey for providing valuable help with the MT-GenEval dataset. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.